OpenStudy (anonymous):

Show all work to solve 1 over x plus 3 over 8=1 over 2. Use complete sentences to explain how to solve this equation.

OpenStudy (anonymous):

1/(x + (3/8)) = 1/2 is the question?

OpenStudy (anonymous):

\[\LARGE \frac1x +\frac38 =\frac12\]

OpenStudy (anonymous):

KRESHNIK IS RIGHT

OpenStudy (anonymous):

redshift I'll leave it to you :)

OpenStudy (anonymous):

1/x + 3/8 = 1/2 So we need to make a common denominator so that we can do fraction addition: We can do this by multiplying each numerator by the other denominator and adding them to make the numerator, and multiply each denominator together to get the common demominator: ( 1(8) + 3(x) )/(8x) = 1/2 then we can cross-multiply: (8 + 3x)(2) = 8x then we expand: 16 + 6x = 8x then we subtract 6x from each side: 16 = 2x then we divide both sides by 2: 8 = x I always hated explaining math in school lol

OpenStudy (anonymous):

\[\LARGE \frac1x=\frac12 -\frac38\] LOL I was somewhere else, when I got back I saw that you were typing ... I thought about doing it ... :P

OpenStudy (anonymous):

I hate my online class :P

OpenStudy (anonymous):

Actually you could do it Kreshnik's way much easier. 1/x = 1/2 - 3/8 Common denominator by multiplying 1 and 2 in 1/2 by 4: 1/x = 4/8 - 3/8 1/x = 1/8 x=9

OpenStudy (anonymous):

x = 8! i mean

OpenStudy (anonymous):

well I don't hate math at all , just because I don't know it, I think people should try harder and not just to say, ahh I hate this ! LOL

OpenStudy (anonymous):

Truee ...:P

OpenStudy (anonymous):

THANKS BTW